# Mantis: Interleaved Multi-Image Instruction Tuning

### v1

@@ Abstract
The recent years have witnessed a great array of large multimodal models (LMMs) to effectively solve single-image vision language tasks. However, their abilities to solve multi-image visual language tasks is yet to be improved. The existing multi-image LMMs (e.g. OpenFlamingo, Emu, Idefics, etc) mostly gain their multi-image ability through pre-training on hundreds of millions of noisy interleaved image-text data from web, which is neither efficient nor effective. In this paper, we aim at building strong multi-image LMMs via instruction tuning with academic-level resources. Therefore, we meticulously construct Mantis-Instruct containing 721K instances from 14 multi-image datasets. We design Mantis-Instruct to cover different multi-image skills like co-reference, reasoning, comparing, temporal understanding. We combine Mantis-Instruct with several single-image visual-language datasets to train our model Mantis to handle any interleaved image-text inputs. We evaluate the trained Mantis on five multi-image benchmarks and eight single-image benchmarks. Though only requiring academic-level resources (i.e. 36 hours on 16xA100-40G), Mantis-8B can achieve state-of-the-art performance on all the multi-image benchmarks and beats the existing best multi-image LMM Idefics2-8B (pre-trained on 140M interleaved image-text data) by an average of 9 absolute points. We observe that Mantis performs equivalently well on the held-in and held-out evaluation benchmarks, which demonstrates its generalization ability. Remarkably, Mantis only lags behind GPT-4V by an average of 2 absolute points across the five benchmarks. We further evaluate Mantis on single-image benchmarks and demonstrate that Mantis can maintain a strong single-image performance on par with CogVLM and Emu2. Our results are particularly encouraging as it shows that low-cost instruction tuning is indeed much more effective than intensive pre-training in terms of building multi-image LMMs. We will release all of our code and data for reproducibility.
Figure 1: Mantis can accept interleaved text-image to perform different tasks like coherence, comparison, logical reasoning, and temporal understanding, etc. Mantis achieves the state-of-the-art performance on the 5 multi-image benchmarks, and preserves abilities for single-image tasks.
@@ 1 Introduction
Large Multimodal Models (LMMs) have advanced significantly in recent years. Both closed-source models like GPT-4V [1], Gemini [2], Reka [3], MM1 [4] and open-source models like LLaVA [5, 6], LLaVA-NeXT [7], BLIP [8], CogVLM [9], Idefics [10] have shown strong visual-language understanding and generation capabilities in single-image tasks like VQA [11], TextVQA [12], GQA [13], etc. Despite the significant interest in improving LMMs’ performance on single-image vision-language tasks, relatively less work attempts to improve LMMs’ capability to solve multi-image vision-language tasks. To the best of our knowledge, only a few open-source models (e.g. Idefics [10], OpenFlamingo [14], Emu2 [15]) are optimized to support multi-image inputs.
@@
We argue that multi-image visual ability is also a crucial ability in real-world applications. Specifically, we categorize the ability into four skills including (1) Co-reference: understanding the references like “second image” in the natural language expression and grounding it on the referred image to generate a response. (2) Reasoning: capturing the information across multiple images and reasoning over these multiple pieces to derive the response. (3) Comparison: capturing the nuances and commonalities between several images. (4) Temporal understanding: observing multiple frames of a video or a scene to understand the temporal information like actions, behaviors, interactions, etc. These are the major four skills we want the LMMs to encompass.
The existing multi-image LMMs like Idefics [10], OpenFlamingo [14], Emu2 [15], MM1 [4] heavily rely on pre-training on a massive amount of interleaved image-text documents like MMC4 [16], Obelics [10] crawled from the web. For example, MMC4 contains 500 million examples, while Obelics contains 140 million examples. Such a pre-training procedure would require a massive amount of computation. Recent works like Mementos [17], BLINK [18], MVBench [19] reveal that they are actually incompetent in handling multi-image inputs despite being extensively pre-trained. We conjecture it’s due to the high volume of noise present in the crawled image-text documents.
In this paper, we attempt to build LMMs with academic-level resources to support interleaved multi-image inputs. Instead of pre-training LMMs with massive amounts of data, we resort to instruction tuning on publicly available academic vision-language datasets. Similar to LLaVA [7], we start from a pre-trained language model like LLaMA-3 [20] or Fuyu [21] and a vision transformer encoder from CLIP [22] or SigLIP [23]. We use the same multimodal projector as LLaVA to map the vision embeddings to the text embeddings space to simply concatenate them with the text embeddings. We also define a text-image interleaving format: “…(image {i}: <BOI>image embeddings</EOI>)…”, where <BOI> and</EOI> are the image delimiters.
@@
We create an instruction-tuning dataset Mantis-Instruct to train our model. Mantis-Instruct has a total of 721K instances, consisting of 14 subsets to cover all the multi-image skills. Among the 14 subsets, 10 subsets are from the existing datasets. For example, NLVR2 [24], IconQA [25], etc are used to cover ‘reasoning’ skill; DreamSim [26], Birds-to-Words [27], etc are used to cover ‘comparing’ skill; NExT-QA [28], STAR [29], etc are used to cover ‘temporal’ skill. We additionally curate four new datasets LLaVA-665k-multi, LRV-multi to cover ‘coref’ skill and Contrast-Caption, Multi-VQA to broaden ‘reasoning’ skill. The detail of all the datasets is included in Table 1. All the subsets are unified in a standardized format, where images and texts are fully interleaved across multi-turn conversations and can be put into any order depending on the task. We train Mantis by combining Mantis-Instruct with another 268K single-image vision-language data to balance the multi-image and single-image abilities.
To evaluate the multi-image skills, we adopt four existing benchmarks NLVR2 [24], QBench [30], BLINK [18], MVBench [19] to cover ‘reasoning’, ‘compare’ and ‘temporal’ skills. We manually create a new challenging dataset Mantis-Eval to cover ‘coref’ and ‘reasoning’ skills. We include 12 competitive baselines including single-image LMMs like LLaVA-Next [7], Qwen-VL [31] and multi-image LMMs like Idefics1/2, OpenFlamingo, and Emu2. For single-image LMMs, we merge multiple images horizontally as a single-image input.
@@
By training on 16 x A100-40G for 36 hours, Mantis achieves the state-of-the-art performance on 5 multi-image tasks, excelling on all the multi-image skills. Mantis is able to outperform the strongest baseline Idefics2 [10] by an average of 9 absolute points. On the held-out evaluation set Mantis-Eval, Mantis beats all the other baselines by at least 9.5 points. These results are encouraging because we show that the low-cost instruction tuning on 721K high-quality data can lead to a much better generalization performance than other LMMs (Emu2, Idefics, etc) intensively pre-trained on 140M-500M noisy web data. Mantis also preserves decent performance of single-image tasks and is on par with CogVLM [9] and Emu2 [15]. We also present insights on the architecture selection and inference approaches based on the results in  section 3. We believe Mantis-Instruct can serve as an essential baseline for future studies on multi-image LMMs due to its efficiency. We will release all the code, data, and models to help the community reproduce our results.
@@ 2.1 Method
LMM architectures are roughly divided into two classes: decoder-only LMM and encoder-decoder LMM. The most popular ones are Fuyu [21] and LLaVA [32]. We modified these two model structure to enable support for interleaved multi-image input. For Fuyu, we preserve the super-resolution feature containing the maximum image size to be 720P. For LLaVA architecture, we also modify the codes to support multi-image inputs.
@@
How many images can a multi-image supported model accept? This is actually limited by both the LLM backbone’s max token length and the number of image tokens the vision encoder generated. For Fuyu, since we limit the image size to 720P, each image consumes at most $10241920\times 1080/30^{2}=$ 1024 image tokens. For the LLaVA architecture, each image consumes a fixed $(336/14)^{2}=576$ image tokens. For Llama3 with maximum window size of 8K, we can feed at most 8 and 14 images respectively. This is fairly enough for all the experiments in this paper.
@@
A proper text-image interleaving format can make an LLM easier to acquire multi-image understanding and reasoning ability. We contend that a good text-image interleaving format should: (1) mark boundaries between images clearly, and (2) denote the serial number of images. Following this principle, we designed our interleaving format as follows: "(image {i}: <BOI><image><EOI>)", where <BOI> is the begin of image token and <EOI> is the end of image token. <image> is the placeholder for image patches. This format add clear separators between images, and gives LMM serial information of the image through "image {i}". In practice, we set <BOI> and <EOI> to be <Image> and </Image> respectively.
Multi-Image Skills:
Training data decides the skills and knowledge that we expect the model to learn. As mentioned in  section 1, 4 crucial multi-image skills, co-reference, reasoning, comparison, and temporal understanding are considered. We also deliberately curate our Mantis-Instruct to cover all these skills. For example, we reformat LLaVA-665k to LLaVA-665k-multi for co-reference skills, and VIST for temporal understanding. Details are stated in the following subsection.
@@ 2.2 Mantis-Instruct: A large scale multiple image question answering dataset
NLVR2 Reason 86K 2.0 2 2.0 105 1257
IconQA Reason 64K 2.4 6 2.0 71 1454
Contrast-Caption Reason 36K 3.8 8 7.6 871 3067
ImageCoDe Reason 17K 2.4 6 1.0 126 5886
Multi-VQA Reason 5K 4.0 6 11.0 1102 3417
@@
We construct Mantis-Instruct from multiple publicly available datasets. Detailed statistics are shown in  Table 1. We use a similar dataset format with LLaVA’s, where each data item contains multiple images and multiple turns of QA pairs are gathered. We report the number of examples, the number of images per item, the average conversation turns, and the average length. We describe the task type of each dataset and the processing techniques used in the followings:
LLaVA-665k-multi: Besides the above-mentioned multi-image reasoning tasks, we also manually reformat the LLaVA-655k dataset from LLaVA-1.5 into a "synthetic" multi-image dataset. We randomly merge 2 to 4 original date items of a single image into a multi-image data item. Then add proper image denotation like "For the second image, …" for each question. Then we shuffle all the conversation pairs to form the final merge items. It’s a "synthetic" multi-image dataset because answering each question in the conversation does not require reasoning across images. However, due to the quality and diversity of the LLaVA-665k data, we contend that it’s helpful for avoiding forgetting via replaying and increasing instruction-following ability.
NLVR2 [24]: NLVR2 is a natural language reasoning dataset grounded in images. The task is to determine whether an statement is true or false given the visual contents of 2 images. NLVR2 covers many linguistic phenomena, such as cardinality, existential quantifiers, universal quantifiers, etc.
ImageCoDe [34]: ImageCoDe is multimodal challenge called image retrieval from contextual descriptions, where a multimodal model is required to retrive/select the correct image from a set of 10 minimally contrastive images. For each image, 9 similar images are retrived from Open Images Dataset V6 [35] using CLIP encodings.
Spot-the-diff [25]: This is a dataset which requires the model to generate the difference description given 2 images. We use the processed version from MIMIC-IT [38].
Multi-VQA: This dataset contains multi-image reasoning QA pairs, where each question is generated by GPT-4 based on the captions of the images. Each question is required to involve at least two images. Images and captions are sampled from ShareGPT-4V [36]. Prompt template is in  Table 5.
Single-image-VQA: Instead of focusing only on multi-image tasks, we claim that it’s necessary to also include some single-image VQA dataset to avoid catastrophic forgetting on single-image tasks. We included some of the single image data from LLaVA-665k as well as DocVQA [41], DVQA [42], and ChartQA [43] to enhance its ability on diagrams and OCR. In practice, since the size of DVQA is too large, we only use 30k output 200k of it for the training.
@@
Figure 2: Illustrations of Mantis-Instruct datasets.
2.3 Training Details
We have trained multiple versions of our models with different dataset-mixing strategies and different backbones. We mainly build three models. Mantis-8B-CLIP and Mantis-8B-SigLIP are based on Llama-3-8B-Instruct [20], where CLIP [22] and SigLIP [23] are used as the visual encoder respectively. Mantis-8B-Fuyu is based on the Adept AI’s Fuyu-8B [21]. Different from the above architecture, there is not visual encoder for the Fuyu model, where each image path is directly mapped to the token embedding space through a linear layer.
For Mantis-8B-CLIP and Mantis-8B-SigLIP, we first pre-train the multimodal projector on LLaVA pre-train data, where learning rate is set to 1e-3, batch size is set to 256. After pre-training, we fine-tune the 2 models on the Mantis-Instruct. For Mantis-8B-Fuyu, we directly fine-tune Fuyu-8B on Mantis-Instruct without pre-training.
During the fine-tuning, we train each model on the data for 1 epoch, with a batch size being 128. The maximum context length is set to 8192. The learning rate is set to 1e-5, which is half of the learning rate used in LLaVA-1.5’s instruction-tuning. We set the warmup ratio to be 0.03 and use a cosine learning rate scheduler. We speed up our training and inference with Flash-attention2 [44].
@@ 3 Experiments
In this section, we aim to evaluate Mantis’s ability on both multi-image and single-image vision-language tasks. For multi-image tasks, depending on whether the LMM can handle multiple images, we either concatenate multiple images (‘merge’) horizontally as a single image or feed multiple images as a ‘sequence’ to the model.
@@ 3.1 Baselines
For ‘merge’ evaluation, we include BLIP-2-13B [8], InstructBLIP-13B [45], Qwen-VL-Chat [31], LLaVA-1.5-7B [5], LLaVA-1.6-7B [7], Kosmos2 [46], and Fuyu-8B [21]. For ‘sequence’ evaluation, we include OpenFlamingo [14], Emu2-Chat-34B [15], Idefics1 [10], Idefics2 [10], GPT-4V [1]. Additionally, for single-image tasks, we include InstructBLIP-7B-Vicuna [45], Yi-VL-6B [47].

### v2

@@ Abstract
Large multimodal models (LMMs) have shown great results in single-image vision language tasks. However, their abilities to solve multi-image visual language tasks is yet to be improved. The existing LMMs like OpenFlamingo, Emu2, Idefics gain their multi-image ability through pre-training on hundreds of millions of noisy interleaved image-text data from the web, which is neither efficient nor effective. In this paper, we aim to build strong multi-image LMMs via instruction tuning with academic-level resources. Therefore, we meticulously construct Mantis-Instruct containing 721K multi-image instruction data to train a family of models Mantis. The instruction tuning empowers Mantis with different multi-image skills like co-reference, comparison, reasoning, and temporal understanding. We evaluate Mantis on five multi-image benchmarks and seven single-image benchmarks. Mantis-SigLIP can achieve SoTA results on all the multi-image benchmarks and beat the strongest multi-image baseline, Idefics2-8B by an average of 11 absolute points. Notably, Idefics2-8B was pre-trained on 140M interleaved multi-image data, which is 200x larger than Mantis-Instruct. We observe that Mantis performs equivalently well on the held-in and held-out benchmarks, which shows its generalization ability. Notably, we found that Mantis can even match the performance of GPT-4V on multi-image benchmarks. We further evaluate Mantis on single-image benchmarks and demonstrate that Mantis also maintains a strong single-image performance on par with CogVLM and Emu2. Our results show that multi-image abilities are not necessarily gained through massive pre-training, instead, it can be gained by the low-cost instruction tuning. Our work provides new perspectives on how to improve LMMs’ multi-image abilities.
Figure 1: Mantis can accept interleaved text-image to perform different multi-image tasks like coherence, comparison, logical reasoning, and temporal understanding. Mantis achieves state-of-the-art performance on the five multi-image benchmarks and preserves abilities for single-image tasks.
@@ 1 Introduction
Large Multimodal Models (LMMs) have advanced significantly in recent years. Both closed-source models like GPT-4V [1], Gemini [2], Reka [3], MM1 [4] and open-source models like LLaVA [5, 6], LLaVA-NeXT [7], BLIP [8], CogVLM [9], Idefics [10] have shown strong visual-language understanding and generation capabilities in single-image tasks like VQA [11], TextVQA [12], GQA [13], etc. Despite the significant interest in improving LMMs’ performance on single-image vision-language tasks, relatively less work attempts to improve LMMs’ capability to solve multi-image vision-language tasks. To the best of our knowledge, only a few open-source models (e.g. Idefics [10], OpenFlamingo [14], Emu2 [15]) and VILA [16] support multi-image inputs.
@@
We argue that multi-image visual ability is also a crucial in real-world applications. Specifically, we categorize the ability into four skills including (1) Co-reference: understanding the references like “second image” in the natural language expression and grounding it on the referred image to generate a response. (2) Comparison: capturing the nuances and commonalities between several images. (3) Reasoning: capturing the information across multiple images and reasoning over these multiple pieces to derive the response. (4) Temporal understanding: observing multiple frames of a video or a scene to understand the temporal information like actions, behaviors, interactions, etc. These are the four major skills we want the LMMs to encompass.
The existing multi-image LMMs like Idefics [10], OpenFlamingo [14], Emu2 [15], MM1 [4], VILA [16] heavily rely on pre-training on a massive interleaved image-text web documents like MMC4 [17], Obelics [10]. For example, MMC4 contains 500 million examples, while Obelics contains 140 million examples. Such a pre-training procedure would require a massive amount of computation, which hampered its adoption in most open-source LMM training pipelines. In this paper, we attempt to build LMMs with academic-level resources to support interleaved multi-image inputs. Specifically, we made a few efforts along this line:
@@
Dataset: We create the first multi-image instruction-tuning dataset Mantis-Instruct. It has a total of 721K instances, consisting of 14 subsets to cover all the multi-image skills. Among the 14 subsets, 10 subsets are from the existing datasets. For example, NLVR2 [18], IconQA [19], etc are used to cover ‘reasoning’ skill; DreamSim [20], Birds-to-Words [21], etc are used to cover ‘comparison’ skill; NExT-QA [22], STAR [23], etc are used to cover ‘temporal understanding’ skill. We additionally curate four new datasets LLaVA-665k-multi, LRV-multi to cover the ‘coref’ skill and Contrast-Caption, Multi-VQA to broaden the ‘reasoning’ skill.
Architecture: Similar to LLaVA [7], we start from a pre-trained language model like LLaMA-3 [24] or Fuyu [25] and a vision transformer encoder from CLIP [26] or SigLIP [27]. We use the same multimodal projector as LLaVA to map the vision embeddings to the text embeddings space to simply concatenate them with the text embeddings. We also define a text-image interleaving format: “…(image {i}: <BOI>image embeddings</EOI>)…”, where <BOI> and</EOI> are the image delimiters. We explored multiple instantiations and found this optimal design.
Evaluation: We manually curate a new challenging dataset Mantis-Eval to cover several multi-image skills. This evaluation can help us better analyze LMMs’ multi-image abilities.
@@
We train Mantis by combining Mantis-Instruct with another 268K single-image vision-language data to balance the multi-image and single-image abilities. To evaluate the multi-image skills, we adopt four existing benchmarks NLVR2 [18], QBench [28], BLINK [29], MVBench [30] and Mantis-Eval to cover all the mentioned skills. We include 15 competitive baseline models including single-image LMMs like LLaVA-Next [7], Qwen-VL [31] and multi-image LMMs like GPT-4V, Idefics1/2, etc. For single-image LMMs, we merge multiple images horizontally as a single-image input. By training on 16 x A100-40G for 36 hours, Mantis achieves state-of-the-art performance on all five multi-image tasks, excelling on all the multi-image skills. Notably, Mantis outperforms the strongest baseline Idefics2 [10] by an average of 11 absolute points. The held-in and held-out results are equivalently strong, which shows the generalization ability of Mantis. These results are encouraging because we show that the low-cost instruction tuning on 721K high-quality data can lead to a much better generalization performance than other LMMs intensively pre-trained on 100x larger datasets. Mantis also preserves its strong performance of single-image tasks and is on par with CogVLM [9] and Emu2 [15]. We also present insights on the architecture selection and inference approaches based on the results in  section 3. We believe Mantis-Instruct can serve as an essential baseline for future studies on multi-image LMMs due to its simplicity and low cost. We will release all the code, data, and models to help with the reproducibility of our results.
@@ 2.1 Method
Most existing LMMs do not support multi-image inputs, either due to the architecture design [32, 8], or the lack of associated code support [33, 25]. To enable multi-image training and inference, we have modified the LLaVA architecture and added code support for it. Idefics2 [34] already supports multi-image as inputs naturally, we thus inherit their structure.
@@
How many images can a multi-image supported model accept? This is limited by both the LLM backbone’s max token length and the number of image tokens the vision encoder generated. For the LLaVA architecture, each image consumes a fixed $(336/14)^{2}=576$ image tokens. For Llama3 with a maximum window size of 8K, we can feed at most 14 images. For Idefics2, the number of tokens per image is resampled to be $64$ tokens only due to the existence of perceiver, which is pretty efficient. Given 8K context length, it can accept 128 images at most, even more than some video models.
@@
A proper text-image interleaving format can make an LLM easier to acquire multi-image understanding and reasoning ability. We contend that a good text-image interleaving format should: (1) mark boundaries between images clearly, and (2) denote the serial number of images. Following this principle, we designed our interleaving format as follows: "(image {i}: <BOI><image><EOI>)", where <BOI> is the begin of image token and <EOI> is the end of image token. <image> is the placeholder for image patches. This format adds clear separators between images, and gives LMM serial information of the image through "image {i}". In practice, we set <BOI> and <EOI> to be <Image> and </Image> respectively. Previous work also demonstrates the effectiveness of this approach through a comprehensive ablation study [35].
@@ 2.2 Mantis-Instruct: A large scale multiple image question answering dataset
NLVR2 Reason 86K 2.0 2 2.0 105 1257
IconQA Reason 64K 2.4 6 2.0 71 1454
Contrast-Caption Reason 36K 3.8 8 7.6 871 3067
ImageCoDe Reason 17K 10 10 1.0 126 5886
Multi-VQA Reason 5K 4.0 6 11.0 1102 3417
@@
We construct Mantis-Instruct from multiple publicly available datasets. Detailed statistics are shown in  Table 1 with several newly curated subset by us. We use a similar dataset format with LLaVA’s, where each data item contains multiple images and multiple turns of QA pairs are gathered. We demonstrate some examples of our newly curated subset in Figure 2. We report the number of examples, the number of images per item, the average conversation turns, and the average length. Mantis-Instruct is collected based on the 4 multi-image skills. We here describe the gathered subsets for each skill briefly. Construction details can be found in  subsection A.1.
Co-reference
requires the model to create a mapping from natural language references, such as "second image," to the actual images in the input. It does not require the model to infer across multiple images. To achieve this, we constructed LLaVA-665k-multi and LRV-multi by concatenating multiple single-image conversations into a multi-image sequence. Deliberately, we included natural language references like "For the second image" for each question.
Comparison
requires the model to not only have good co-reference ability but also be able to compare and understand differences across multiple images. We gather Co-Instruct, Dreamsim, Spot-the-Diff, and Birds-to-Words together to enhance the model’s comparison ability. They cover a wide range of topics such as image quality, visual similarity, and difference description.
Reasoning
requires the model to further make inferences by combining its world knowledge with the information collected using the co-reference and comparison ability. It distinguishes from the comparison skill by introducing more complex human instruction associated with the images. We gather NLVR2, IconQA, Contrast-Caption by reformatting captioning datasets, ImageCoDe, and our self-collected Multi-VQA to further equip the model with reasoning ability. The topics include logical reasoning, counting, image matching, image retrieval, and free-form multi-image QA.
Temporal Understanding
requires the model to understand temporal image sequences such as videos, comics, etc. We include a small number of video understanding datasets (14K), VIST, NExT-QA, and STAR, to activate the model’s temporal understanding ability. We contend that a model with the above 3 multi-image skills shall be easily tuned to understand image sequences.
@@
Figure 2: Illustrations of Mantis-Instruct datasets from Contrast-Captioning and Multi-VQA.
@@ 3 Experiments
In this section, we aim to evaluate Mantis’s ability on both multi-image and single-image vision-language tasks. For multi-image tasks, depending on whether the LMM can handle multiple images, we either concatenate multiple images (‘merge’) horizontally as a single image or feed multiple images as a ‘sequence’ to the model. We have trained four model variants listed in  Table 2, where we list the architecture and dataset details. Mantis-CLIP and Mantis-SigLIP only adopt the CC3M 560K subset for feature alignment. Mantis-Flamingo and Mantis-Idefics2 are initialized from OpenFlamingo and Idefics2, respectively, which have been intensively pre-trained on the multi-image corpus. After this pre-training and preparation, we fine-tune them on Mantis-Instruct to get Mantis. Please refer to  subsection A.3 for details of hyper-parameters and settings.
Table 2: Architecture and training details of mantis model family. "I" for "Image" in last column.
Mantis-Models V-Encoder LLM-backbone Pre-Training (PT) PT size Fine-Tuning # Tokens / I
Mantis-CLIP CLIP LLaMA-3-8B CC3M subset 0.56M Mantis-Instruct 576
Mantis-SIGLIP SigLIP LLaMA-3-8B CC3M subset 0.56M Mantis-Instruct 576
Mantis-Flamingo CLIP MPT-7B MMC4, LAION 2421M Mantis-Instruct 576
Mantis-Idefics2 SiGLIP Mistral-7B-v0.1 OBELICS, Cauldron 143M Mantis-Instruct 64
@@ 3.1 Baselines
For ‘merge’ evaluation, we include BLIP-2-13B [8], InstructBLIP-13B [36], Qwen-VL-Chat [31], LLaVA-1.5-7B [5], LLaVA-1.6-7B [7], Kosmos2 [33], and Fuyu-8B [25]. For ‘sequence’ evaluation, we include OpenFlamingo [14], Otter [37], VideoLLaVA [38], Emu2-Chat-34B [15], VILA [39], Idefics1 [10], Idefics2 [10], GPT-4V [1]. Additionally, for single-image tasks, we include InstructBLIP-7B-Vicuna [36], Yi-VL-6B [40].